OpenStudy (anonymous):

A 10-V battery is connected to an RC circuit (R = 6 Ω and C = 10 μF). Initially, the capacitor is uncharged. What is the final charge on the capacitor (in μC)?

OpenStudy (jamesj):

Clearly in equilibrium the potential difference of the capacitor is 10 volts. Now just use the definition of capacitance, C = Q/V to calculate Q.

OpenStudy (anonymous):

so the final charge would be 100?

OpenStudy (rulnick):

Final charge is 100 microcoulombs. I answered in math for you. See work there.

OpenStudy (anonymous):

could you help with this one? The quantity t1/2=τ ln 2 is called the half-life of an exponential decay, where τ=RC is the time constant in an RC circuit. The current in a discharging RC circuit drops by half whenever t increases by t1/2. For a circuit with R=1 kΩ and C=4 μF, if the current is 6 mA at t=4 ms, at what time (in ms) will the current be 3 mA?

OpenStudy (radar):

For "practical" matters you may consider the capacitor fully charged after 5 time constants.

OpenStudy (radar):

Where the time constant is equal to RC.